ROTATIONAL SPECTRA OF THE Kr-H_{2}O VAN DER WAALS COMPLEX

\underline{JENNIFER VAN WIJNGAARDEN} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, Canada, T6G 2G2.

Rotational spectra of the Kr-H_{2}O van der Waals complex were measured between 4 and 19 GHz using a pulsed jet cavity Fourier transform microwave spectrometer. The isotopomers studied include those of H_{2}O, HDO, D_{2}O, H_{2}^{17}O, and H_{2}^{18}O with 6 different isotopes of Kr. Tunnelling splittings were observed for all isotopomers with the exception of Kr-HDO. Transitions of complexes with ^{83}Kr, D_{2}O, HDO, and H_{2}^{17}O showed nuclear hyperfine structure due to the quadrupole moments of ^{83}Kr, D, and ^{17}O, respectively. Rotational and centrifugal distortion constants were determined and used to calculate structural parameters.